# Learning DFAs from Positive Examples Only via Word Counting

### v1

@@ 1 Introduction
In recent years, the rapid advancement and proliferation of Artificial Intelligence systems have transformed numerous fields, from healthcare zhang2022applications to transportation bharadiya2023artificial or finance chen2023explainable. This surge in AI capabilities has greatly increased the need of interpretable models able to provide insights into the decision-making processes of complex black-box systems.
@@
Interpretable models are synthesized from temporal data. The goal is to obtain a model that is coherent with the observations gathered, which in turn can be used to decipher the system’s past behavior. Different kinds of (interpretable) models are used in the literature, often of one of two types: finite-state machines weiss2024extracting and (temporal) logic formulas DBLP:conf/aips/CamachoM19. In this paper, we focus on Deterministic Finite Automata (DFAs) DBLP:journals/ibmrd/RabinS59, that enjoy both an accessible formalism, making them suitable e.g. classifier of sequential data DBLP:conf/aaai/ShvoLIM21, and many desirable properties, such as tractable model checking.
@@
In the literature, synthesizing a DFA from observations is usually formalized as a passive learning task where we are given a positive set of words that a prospective DFA should accept, a negative set of words that it should reject, and an integer bound on its number of states DBLP:journals/iandc/Gold78; DBLP:conf/cade/GrinchteinLP06; DBLP:conf/icgi/HeuleV10. However, a significant obstacle on the usefulness of the passive learning task from positive and negative examples (PLPN) is the difficulty of coming up with negative examples: they correspond to observations of a system’s undesirable behaviors, which can be intrinsically hard (e.g., with black-box systems) or unrealistic (e.g., with self-driving cars) to gather.
@@
To circumvent that issue, the focus has recently shifted to the passive learning task of synthesizing DFAs from positive examples only (PLP). Without any additional restrictions, the DFA accepting exactly the (positive) words of the input set $\mathcal{P}$—on the one end of the spectrum—and the trivial accepting DFA (the one-state DFA accepting every word)—on the other end of the spectrum—both meet the requirement of this PLP task while providing no insight whatsoever w.r.t. the input set $\mathcal{P}$. To forbid these two extreme solutions, constraints on the prospective DFA are added. On the one hand, as in the classical PLPN setting, we consider a bound $n\in\mathbb{N}$ on the number of states of the prospective DFA. The benefit is twofold: first, it prevents synthesizing the DFA accepting exactly the words in $\mathcal{P}$ (assuming $n$ is small enough); second, for explainability purposes, the smaller the DFA, the better. On the other hand, we look for a DFA that is language-minimal, i.e., whose language is minimal (for inclusion) among those DFAs with at most $n$ states that accept all words in $\mathcal{P}$. As argued in DBLP:conf/icgi/AvellanedaP18, this constraint follows naturally from a parsimony standpoint: we look for a simplest solution. With these restrictions, we obtain a PLP task $\mathcal{T}$.
@@
By construction, the symbolic algorithm avoids redundancies that could occur in the CE algorithm. However, even with the symbolic algorithm, a significant drawback remains: the best bound on the number of steps of the algorithm, and thus on the number of calls to SAT solvers, is exponential in $n$. This is due to the fact that there are exponentially many words of length at most (polynomial in) $n$. Furthermore, while the complexity of the PLPN task (which can be stated as a decision problem) is well understood—it is known to be NP-complete since the 70s DBLP:journals/iandc/Gold78—the PLP task is poorly understood complexity-wise. In particular, DBLP:conf/icgi/AvellanedaP18; DBLP:conf/aaai/0002GBN0T23 do not provide a complexity lower bound on any closely-related decision problem. As such, it is not clear that solving the PLP task $\mathcal{T}$ even requires the use of SAT solvers.
@@
In this paper, we tackle the PLP task $\mathcal{T}$ from a new perspective which consists in assigning to each DFA a word-counting metric that greatly enhances our ability to compare two DFAs (as opposed to checking language inclusion). Specifically, this word-counting metric is the number of words accepted by each DFA up to length 2n22n-2. To demonstrate the usefulness of this new perspective, we show that a word-counting minimal DFA is necessarily language-minimal [^1] The length 2n22n-2 is optimal in that regard (i.e., a smaller length would not satisfy this property).. Thus, we focus on a new PLP task $\mathcal{T}^{\prime}$—that requires a prospective DFA to be word-counting minimal instead of only language-minimal—and on the associated decision problem about the minimal word-counting metric of DFAs with at most $n$ states accepting all words in $\mathcal{P}$. We show that this problem is $\mathsf{NP}$-complete, which constitutes the first complexity result directly related to a PLP task. This $\mathsf{NP}$-completeness result has two main consequences: first, because it is in $\mathsf{NP}$, we immediately obtain that the task $\mathcal{T}^{\prime}$, and thus the task $\mathcal{T}$ as well, can be solved with polynomially many calls to a SAT solver, via a binary search; second, because the problem is $\mathsf{NP}$-hard, it is unlikely that we can do better than polynomially many calls to a SAT solver to solve task $\mathcal{T}^{\prime}$ (though task $\mathcal{T}$ could be intrinsically easier to solve than task $\mathcal{T}^{\prime}$).
@@
As already mentioned, there is a large body of work focusing the classical DFA passive learning. Another widely studied DFA learning setting is active learning, where a learner successively queries a teacher. Angluin’s seminal work DBLP:journals/iandc/Angluin87 essentially started the research on this topic, and is still very influential nowadays DBLP:conf/fm/ShahbazG09; DBLP:conf/ijcai/BolligHKL09.
DFA learning from positive examples only was introduced in DBLP:journals/iandc/Gold67, where it is shown that DFAs cannot be identified at the limit. This partly explains why there was no further study of this topic until recently DBLP:conf/icgi/AvellanedaP18; DBLP:conf/aaai/0002GBN0T23. More popular subjects are the passive learning from positive examples only of better-behaved models, such as pattern languages DBLP:journals/jcss/Angluin80 hidden Markov chains stolcke1992hidden, or stochastic machines DBLP:journals/ita/CarrascoO99.
Finally, an ILP encoding a learning task (of a logical (LTLf) formula) was first introduced in DBLP:conf/ilp/IeloLFRGR23.
Many technical details are postponed to the Appendix.
@@ 2 Definitions and notations
Given an integer $n\in\mathbb{N}$ and a set of words $\mathcal{P}$, we let $\mathsf{Rec}(\mathcal{P},n):=\{\mathcal{A}\in\mathsf{DFA}(\Sigma,n)\mid\mathcal{P}\subseteq\mathcal{L}(\mathcal{A})\}$ denote the set of DFAs with at most $n$ states that accept all words in $\mathcal{P}$.
Partial orders.
A relation $\prec\subseteq\mathsf{DFA}(\Sigma)\times\mathsf{DFA}(\Sigma)$ is a partial order on $\mathsf{DFA}(\Sigma)$ if it is reflexive, antisymmetric, and transitive. For $S\subseteq\mathsf{DFA}(\Sigma)$ and $\mathcal{A}\in S$, $\mathcal{A}$ is said to be $S$-minimal w.r.t. $\prec$ if, for all $\mathcal{A}^{\prime}\in S$, we do not have $\mathcal{A}^{\prime}\prec\mathcal{A}$.
Given two DFAs $\mathcal{A},\mathcal{A}^{\prime}\in\mathsf{DFA}$, we write $\mathcal{A}\prec_{\mathcal{L}}\mathcal{A}^{\prime}$ when $\mathcal{L}(\mathcal{A})\subsetneq\mathcal{L}(\mathcal{A}^{\prime})$. Clearly, $\prec_{\mathcal{L}}$ is a partial order on $\mathsf{DFA}(\Sigma)$.
@@ 3 Learning DFAs via word counting
Before we detail our appraoch for solving Task 2, let us describe the state-of-the-art (symbolic) algorithm.
@@
It was developed by Roy et al. (2023) and proceeds as follows. Initially, one starts with the trivial accepting DFA $\mathcal{A}_{0}\in\mathsf{Rec}(\mathcal{P},n)$. Then, at step $i\geq 1$, one computes a DFA $\mathcal{A}_{i}\in\mathsf{Rec}(\mathcal{P},n)$ satisfying $\mathcal{A}_{i}\prec_{\mathcal{L}}\mathcal{A}_{i-1}$ using a SAT solver. The process stops when no such DFA is found. Overall, this algorithm iteratively computes a decreasing chain of DFAs $\mathcal{A}_{0}\succ_{\mathcal{L}}\ldots\succ_{\mathcal{L}}\mathcal{A}_{k}\succ_{\mathcal{L}}\ldots$ with $\mathcal{A}_{i}\in\mathsf{Rec}(\mathcal{P},n)$, for all $i\in\mathbb{N}$. Since the languages of DFAs may be infinite, it may not be clear that this chain is necessarily finite. However, it is indeed the case, as argued in DBLP:conf/aaai/0002GBN0T23: two DFAs with at most $n$ states have the same language if they accept exactly the same words of length at most $n^{2}$ (a tighter bound exists, as discussed below). Thus, since there are finitely many words of length at most $n^{2}$, the above decreasing chain of DFAs is necessarily finite. However, the best bound on the number of iterations that we can extract from this argument is exponential in $n$ (since there are exponentially many words of length at most $n^{2}$), with each iteration involving a call to a SAT solver.
@@
To design a method that does significantly less calls to SAT solvers, we shift perspective from focusing on language inclusion, via the partial order $\prec_{\mathcal{L}}$, to focusing on a numerical order induced by values that we assign to DFAs. As we will see in the remainder of this paper, this has two main benefits: first, numerical orders are compatible with binary searches; second, the values assigned to DFAs can be seen as a score that we aim at optimizing (in this case, minimize). However, it is crucial that this numerical order relates to the language-inclusion order $\prec_{\mathcal{L}}$ previously defined, as this is the one involved in Task 2. In fact, taking as values assigned to DFAs the number of words, up to a certain length, that they accept does the job. This is formally defined below.
@@
In other words, we have $\mathcal{A}\prec_{h}\mathcal{A}^{\prime}$ when $\mathcal{A}$ accepts fewer words of length at most $h$ than $\mathcal{A}^{\prime}$. Clearly, for all $\in\mathbb{N}, h\prec_{h}$ is a partial order on DFAs. Furthermore, this order satisfies the following property, thus showing that minimizing the number of accepted words also minimizes the language.
@@
This proposition relies of the folk theorem recalled below.
@@
We provide a proof of this theorem in Appendix A. It essentially relies on the iterative computation of the Myhill-Nerode equivalence classes of DFAs. In that same section, we also show that the bound $h=|\mathcal{A}|+|\mathcal{A}^{\prime}|-2$ is tight.
@@
Note that this lemma would not hold with automata that are not deterministic. The (simple) proof of this lemma can be found in Appendix B.
@@
Second, since Decision Problem 7 is $\mathsf{NP}$-hard, unless $\mathsf{P}=\mathsf{NP}$, it seems unlikely that it is possible to do, asymptotically, significantly better than the bound of Proposition 10 to solve Task 6. As this only applies to Task 6, thus there might be an efficient way to solve Task 2 without solving Task 6.
@@
The NP-hardness proof of Theorem 9 is quite technical. The formal definition of the reduction, along with the careful proof of its correctness can be found in Appendix B Here, we only give a bird’s eye view of the steps that we take to prove the result.
@@
For the remainder of this section, we consider an instance $(X,\mathcal{C})$ of Problem 11 with $X=\{x_{i}\mid i\in\llbracket 1,r\rrbracket\}$ the set of variables, and $\mathcal{C}=\{C_{j}\mid j\in\llbracket 1,s\rrbracket\}=\mathcal{C}_{+}\uplus\mathcal{C}_{-}$ the set of clauses. The alphabet that we consider is $\Sigma:=\{a,b\}$.
Our goal is to define a set of positive words $\mathcal{P}$ and two integers $n$ and $k$ such that $(X,\mathcal{C})$ is a positive instance of Problem 11 if and only if there is a $(\mathcal{P},n,k)$-suitable DFA $\mathcal{A}$, i.e., such that 1) $\mathcal{A}\in\mathsf{Rec}(\mathcal{P},n)$; and 2) $|\Sigma^{\leq 2n-2}\cap(\mathcal{L}(\mathcal{A})\setminus\mathcal{P})|\leq k$. Condition 2) refers to the number of words of length at most 2n22n-2 not in $\mathcal{P}$ accepted by the DFA, which is different from the setting of Problem 7. This change is harmless—up to considering, as input of Problem 7, $k^{\prime}:=k+|\Sigma^{\leq 2n-2}\cap\mathcal{P}|$. We shift perspective (in this section only), because the words that are accepted that should not because are easier to reason about.
@@
Consider the set of variables $X=\{x_{1},x_{2},x_{3}\}$, a single positive clause $C_{1}=\{x_{1},x_{3}\}\in\mathcal{C}_{+}$ and single a negative clause $C_{2}=\{x_{2},x_{3}\}\in\mathcal{C}_{-}$. In Figure 1, we scheme the shape of a DFA $\mathcal{A}_{\mathsf{ex}}$ that would be $(\mathcal{P},n,k)$-suitable, with $(\mathcal{P},n,k)$ being the result of our reduction from the instance $(X,\mathcal{C})$ of Problem 11. The double-circled states are the accepting states.
In $\mathcal{P}$, we define a $\top$-word” $a\cdot a\cdot a$ and a $\bot$-word” $a\cdot b\cdot a\cdot a$. One can see that when the DFA $\mathcal{A}_{\mathsf{ex}}$ runs on those words, it visits the right-most (green and red) states.
@@
Let us now describe a little more formally the steps that we take to ensure that $(X,\mathcal{C})$ is a positive instance of Problem 11 if and only if there is a $(\mathcal{P},n,k)$-suitable DFA $\mathcal{A}$.
@@
With Items 5. and 6., we have that the valuation defined in Item 3. satisfies $(X,\mathcal{C})$. Indeed, consider e.g. a positive clause $C\in\mathcal{C}$. By Fact 5, there is variable $x_{C}\in C$ such that $\delta^{*}(q_{0},w_{C}\cdot w_{\mathsf{Cl}})=\delta^{*}(q_{0},w_{x_{C}})$. By Fact 6, we have $\delta^{*}(q_{0},w_{x_{C}}\cdot w_{\mathsf{Var}})=\delta^{*}(q_{0},w_{C}\cdot w_{\mathsf{Cl}}\cdot w_{\mathsf{Var}})=\delta^{*}(q_{0},w_{\top})$. Thus, $\nu(x_{C})=\top$ and the valuation $\nu$ satisfies the clause $C$.
In our actual reduction—described in Appendix B—there are many words in $\mathcal{P}$ associated with $\top$, $\bot$, variables and clauses. In that section, we draw (almost) entirely a $(\mathcal{P},n,k)$-suitable DFA built from our reduction on the input of Problem 11 from Examples 12. That DFA has much more states that the one depicted in Figure 1 (which is why it is not depicted here).

### v2

@@ 1 Introduction
In recent years, the rapid advancement and proliferation of Artificial Intelligence systems have transformed numerous fields, from healthcare [zhang2022applications] to transportation [bharadiya2023artificial] or finance [chen2023explainable]. This surge in AI capabilities has greatly increased the need of interpretable models able to provide insights into the decision-making processes of complex black-box systems.
@@
Interpretable models are synthesized from temporal data. The goal is to obtain a model that is coherent with the observations gathered, which in turn can be used to decipher the system’s past behavior. Different kinds of (interpretable) models are used in the literature, often of one of two types: finite-state machines [weiss2024extracting] and (temporal) logic formulas [DBLP:conf/aips/CamachoM19]. In this paper, we focus on Deterministic Finite Automata (DFAs) [DBLP:journals/ibmrd/RabinS59], that enjoy both an accessible formalism, making them suitable e.g. classifier of sequential data [DBLP:conf/aaai/ShvoLIM21], and many desirable properties, such as tractable model checking.
@@
In the literature, synthesizing a DFA from observations is usually formalized as a passive learning task where we are given a positive set of words that a prospective DFA should accept, a negative set of words that it should reject, and an integer bound on its number of states [DBLP:journals/iandc/Gold78, DBLP:conf/cade/GrinchteinLP06, DBLP:conf/icgi/HeuleV10]. However, a significant obstacle on the usefulness of the passive learning task from positive and negative examples (PLPN) is the difficulty of coming up with negative examples: they correspond to observations of a system’s undesirable behaviors, which can be intrinsically hard (e.g., with black-box systems) or unrealistic (e.g., with self-driving cars) to gather.
@@
To circumvent that issue, the focus has recently shifted to the passive learning task of synthesizing DFAs from positive examples only (PLP). Without any additional restriction, the DFA accepting exactly the (positive) words of the input set $\mathcal{P}$—on the one end of the spectrum—and the trivial accepting DFA (the one-state DFA accepting every word)—on the other end of the spectrum—both meet the requirement of this PLP task while providing no insight whatsoever w.r.t. the input set $\mathcal{P}$. To forbid these two extreme solutions, constraints on the prospective DFA are added. On the one hand, as in the classical PLPN setting, we consider a bound $n\in\mathbb{N}$ on the number of states of the prospective DFA. The benefit is twofold: first, it prevents synthesizing the DFA accepting exactly the words in $\mathcal{P}$ (assuming $n$ is small enough); second, for explainability purposes, the smaller the DFA, the better. On the other hand, we look for a DFA that is language-minimal, i.e., whose language is minimal (for inclusion) among those DFAs with at most $n$ states that accept all words in $\mathcal{P}$. As argued by Avellaneda and Petrenko (2018), this constraint follows naturally from a parsimony standpoint: we look for a simplest solution. With these restrictions, we obtain a PLP task $\mathcal{T}$.
@@
By construction, the symbolic algorithm avoids redundancies that could occur in the CE algorithm. However, even with the symbolic algorithm, a significant drawback remains: the best bound on the number of steps of the algorithm, and thus on the number of calls to SAT solvers, is exponential in $n$. This is due to the fact that there are exponentially many words of length at most (polynomial in) $n$. Furthermore, while the complexity of the PLPN task (which can be stated as a decision problem) is well understood—it is known to be NP-complete since the 70s [DBLP:journals/iandc/Gold78]—the PLP task is poorly understood complexity-wise. In particular, [DBLP:conf/icgi/AvellanedaP18, DBLP:conf/aaai/0002GBN0T23] do not provide a complexity lower bound on any closely-related decision problem. As such, it is not clear that solving the PLP task $\mathcal{T}$ even requires the use of SAT solvers.
@@
In this paper, we tackle the PLP task $\mathcal{T}$ from a new perspective which consists in assigning to each DFA a word-counting metric that greatly enhances our ability to compare two DFAs (as opposed to checking language inclusion). Specifically, this word-counting metric is the number of words accepted by each DFA up to length 2n22n-2. To demonstrate the usefulness of this new perspective, we show that a word-counting minimal DFA is necessarily language-minimal. Thus, we focus on a new PLP task $\mathcal{T}^{\prime}$—that requires a prospective DFA to be word-counting minimal instead of only language-minimal—and on the associated decision problem about the minimal word-counting metric of DFAs with at most $n$ states accepting all words in $\mathcal{P}$. Our main contribution is that this problem is $\mathsf{NP}$-complete, which constitutes the first complexity result directly related to a PLP task. This $\mathsf{NP}$-completeness result has two main consequences: first, because it is in $\mathsf{NP}$, we immediately obtain that the task $\mathcal{T}^{\prime}$, and thus the task $\mathcal{T}$ as well, can be solved with polynomially many calls to a SAT solver, via a binary search; second, because the problem is $\mathsf{NP}$-hard, it is unlikely that we can do better than polynomially many calls to a SAT solver to solve task $\mathcal{T}^{\prime}$ (though task $\mathcal{T}$ could be intrinsically easier to solve than task $\mathcal{T}^{\prime}$).
@@
Missing technical details can be found in the extended version [bordais2025learningdfaspositiveexamples].
@@
As already mentioned, there is a large body of work focusing on the classical DFA passive learning. Another widely studied DFA learning setting is active learning, where a learner successively queries a teacher. Angluin’s seminal work [DBLP:journals/iandc/Angluin87] essentially started the research on this topic, and is still very influential nowadays [DBLP:conf/fm/ShahbazG09, DBLP:conf/ijcai/BolligHKL09].
DFA learning from positive examples only was introduced by Gold (1967), where it is shown that DFAs cannot be identified at the limit. This partly explains why there was no further study of this topic until recently [DBLP:conf/icgi/AvellanedaP18, DBLP:conf/aaai/0002GBN0T23]. More popular subjects are the passive learning from positive examples only of better-behaved models, such as pattern languages [DBLP:journals/jcss/Angluin80], hidden Markov chains [stolcke1992hidden], or stochastic machines [DBLP:journals/ita/CarrascoO99].
Finally, an ILP encoding a learning task (of a logical (LTLf) formula) was first introduced in [DBLP:conf/ilp/IeloLFRGR23].
@@ 2 Definitions and notations
Given an integer $n\in\mathbb{N}$ and a set of words $\mathcal{P}$, we let $\mathsf{Rec}(\mathcal{P},n):=\{\mathcal{A}\in\mathsf{DFA}(\Sigma)\mid|\mathcal{A}|\leq n,\;\mathcal{P}\subseteq\mathcal{L}(\mathcal{A})\}$ denote the set of DFAs with at most $n$ states that accept all words in $\mathcal{P}$.
Strict partial orders.
A relation $\prec\subseteq\mathsf{DFA}(\Sigma)\times\mathsf{DFA}(\Sigma)$ is a strict partial order on $\mathsf{DFA}(\Sigma)$ if it is irreflexive, antisymmetric, and transitive. For $S\subseteq\mathsf{DFA}(\Sigma)$ and $\mathcal{A}\in S$, $\mathcal{A}$ is said to be $S$-minimal w.r.t. $\prec$ if, for all $\mathcal{A}^{\prime}\in S$, we do not have $\mathcal{A}^{\prime}\prec\mathcal{A}$.
Given two DFAs $\mathcal{A},\mathcal{A}^{\prime}\in\mathsf{DFA}$, we write $\mathcal{A}\prec_{\mathcal{L}}\mathcal{A}^{\prime}$ when $\mathcal{L}(\mathcal{A})\subsetneq\mathcal{L}(\mathcal{A}^{\prime})$. Clearly, $\prec_{\mathcal{L}}$ is a strict partial order on $\mathsf{DFA}(\Sigma)$.
@@ 3 Learning DFAs via word counting
Before we detail our approach for solving Task 2, let us describe the state-of-the-art (symbolic) algorithm.
@@
It was developed by Roy et al. (2023) and proceeds as follows. Initially, one starts with the trivial accepting DFA $\mathcal{A}_{0}\in\mathsf{Rec}(\mathcal{P},n)$. Then, at step $i\geq 1$, one computes a DFA $\mathcal{A}_{i}\in\mathsf{Rec}(\mathcal{P},n)$ satisfying $\mathcal{A}_{i}\prec_{\mathcal{L}}\mathcal{A}_{i-1}$ using a SAT solver. The process stops when no such DFA is found. Overall, this algorithm iteratively computes a decreasing chain of DFAs $\mathcal{A}_{0}\succ_{\mathcal{L}}\ldots\succ_{\mathcal{L}}\mathcal{A}_{k}\succ_{\mathcal{L}}\ldots$ with $\mathcal{A}_{i}\in\mathsf{Rec}(\mathcal{P},n)$, for all $i\in\mathbb{N}$. Since the languages of DFAs may be infinite, it may not be clear that this chain is necessarily finite. However, it is indeed the case, as argued by Roy et al. (2023): two DFAs with at most $n$ states have the same language if they accept exactly the same words of length at most $n^{2}$ (a tighter bound exists, as discussed below). Thus, since there are finitely many words of length at most $n^{2}$, the above decreasing chain of DFAs is necessarily finite. However, the best bound on the number of iterations that we can extract from this argument is exponential in $n$ (since there are exponentially many words of length at most $n^{2}$), with each iteration involving a call to a SAT solver.
@@
To design a method that does significantly less calls to SAT solvers, we shift perspective from focusing on language inclusion, via the strict partial order $\prec_{\mathcal{L}}$, to focusing on a numerical order induced by values that we assign to DFAs. As we will see in the remainder of this paper, this has two main benefits: first, numerical orders are compatible with binary searches; second, the values assigned to DFAs can be seen as a score that we aim at optimizing (in this case, minimize). However, it is crucial that this numerical order relates to the language-inclusion order $\prec_{\mathcal{L}}$ previously defined, as this is the one involved in Task 2. In fact, taking as values assigned to DFAs the number of words, up to a certain length, that they accept does the job. This is formally defined below.
@@
In other words, we have $\mathcal{A}\prec_{h}\mathcal{A}^{\prime}$ when $\mathcal{A}$ accepts fewer words of length at most $h$ than $\mathcal{A}^{\prime}$. Clearly, for all $\in\mathbb{N}, h\prec_{h}$ is a strict partial order on DFAs. Furthermore, this order satisfies the following property, thus showing that minimizing the number of accepted words also minimizes the language.
@@
This proposition relies on the folk theorem recalled below.
@@
We provide a proof of this theorem in [bordais2025learningdfaspositiveexamples]. It essentially relies on the iterative computation of the Myhill-Nerode equivalence classes of DFAs. We also show that the bound $h=|\mathcal{A}|+|\mathcal{A}^{\prime}|-2$ is tight.
@@
Note that this lemma would not hold with automata that are not deterministic.
@@
Second, unless $\mathsf{P}=\mathsf{NP}$, the $\mathsf{NP}$-hardness of Decision Problem 7 makes it unlikely that it is possible to solve Task 6 with an asymptotic bound significantly better than that of Proposition 10. This only applies to Task 6, and there might be an efficient way to solve Task 2 without solving Task 6.
@@
The NP-hardness proof of Theorem 9 is quite technical, see [bordais2025learningdfaspositiveexamples] for a detailed proof. Here, we only give a bird’s eye view of the steps that we take to prove the result.
@@
For the remainder of this section, we consider an instance $(X,\mathcal{C})$ of Problem 11 with $X=\{x_{i}\mid i\in\llbracket 1,r\rrbracket\}$ the set of variables, and $\mathcal{C}=\{C_{j}\mid j\in\llbracket 1,s\rrbracket\}=\mathcal{C}_{+}\uplus\mathcal{C}_{-}$ the set of clauses. The alphabet that we consider is $\Sigma:=\{a,b\}$. Our goal is to define a set of positive words $\mathcal{P}$ and two integers $n$ and $k$ such that $(X,\mathcal{C})$ is a positive instance of Problem 11 if and only if there is a $(\mathcal{P},n,k)$-suitable DFA $\mathcal{A}$, i.e., such that 1) $\mathcal{A}\in\mathsf{Rec}(\mathcal{P},n)$; and 2) $|\Sigma^{\leq 2n-2}\cap(\mathcal{L}(\mathcal{A})\setminus\mathcal{P})|\leq k$ [^1] This condition constitutes a slight change compared to Problem 7. This is however harmless, see the extended version..
@@
Consider the set of variables $X=\{x_{1},x_{2},x_{3}\}$, a single positive clause $C_{1}=\{x_{1},x_{3}\}\in\mathcal{C}_{+}$ and single a negative clause $C_{2}=\{x_{2},x_{3}\}\in\mathcal{C}_{-}$. In Figure 1, we scheme the shape of a DFA $\mathcal{A}_{\mathsf{ex}}$ that would be $(\mathcal{P},n,k)$-suitable, with $(\mathcal{P},n,k)$—which we partly describe here—being the result of the reduction from the instance $(X,\mathcal{C})$. The double-circled states are accepting.
In $\mathcal{P}$, we define a $\top$-word” $a\cdot a\cdot a\cdot a$ and a $\bot$-word” $a\cdot b\cdot a\cdot a$. One can see that when the DFA $\mathcal{A}_{\mathsf{ex}}$ runs on those words, it visits the right-most (green and red) states.
@@
Let us now describe a little more formally how the words that we define in $\mathcal{P}$ ensure that the existence of a $(\mathcal{P},n,k)$-suitable DFA $\mathcal{A}$ implies that $(X,\mathcal{C})$ is a positive instance of Problem 11. Below, we consider a DFA $\mathcal{A}=(Q,\Sigma,q_{\mathsf{init}},\delta,F)$ satisfying conditions 1) and 2).
@@
With Items 5. and 6., we have that the valuation defined in Item 3. satisfies $(X,\mathcal{C})$. Indeed, consider e.g. a positive clause $C\in\mathcal{C}$. By Fact 5, there is variable $x_{C}\in C$ such that $\delta^{*}(q_{0},w_{C}\cdot w_{\mathsf{Cl}})=\delta^{*}(q_{0},w_{x_{C}})$. By Fact 6, we have $\delta^{*}(q_{0},w_{x_{C}}\cdot w_{\mathsf{Var}})=\delta^{*}(q_{0},w_{C}\cdot w_{\mathsf{Cl}}\cdot w_{\mathsf{Var}})=\delta^{*}(q_{0},w_{\top})$. Thus, $\nu(x_{C})=\top$ and the valuation $\nu$ satisfies the clause $C$. Note that, in the actual reduction described in the extended version [bordais2025learningdfaspositiveexamples], there are many words in $\mathcal{P}$ associated with $\top$, $\bot$, variables and clauses.